In a local market, the monthly price of Internet access service decreases from $40 to $30, and the total quantity of monthly accounts across all Internet access providers increases from 100,000 to 200,000.

What is the value price elasticity of demand, expressed as a positive number? ____. (Round your answer to two decimal places.)
Answer #1

Answer

Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=200000-100000=100000
average quantity=(200000+100000)/2=150000
change in price=30-40=-10
average price=(30+40)/2=35
Elasticity of demand=(100000/150000)/(-10/35)

=-2.33333333

=-2.33

=2.33 (abolute value)

the value price elasticity of demand, expressed as a positive number is 2.33.